Find the angle of elevation to the top of a 56 m high building from point A which is 113 m

from its base. What is the angle of depression from the top of the building to A? /6mks

(b) The expression 6x^2+ x +7 leaves the same remainder when divided by x – m and by x+2m,


where m ≠ 0. Calculate the value of m./4mks


1
Expert's answer
2021-09-06T13:08:22-0400

(a) Let "\theta=" the angle of elevation to the top of a building from point "A."

Consider right triangle "BAC"


"\tan \theta=\dfrac{opposite}{adjacent}=\dfrac{BC}{AC}=\dfrac{56 m}{113 m}"

"\theta=\tan^{-1}\dfrac{56}{113}\approx26.36\degree"

Let "\alpha=" the angle of depression from the top of the building to "A."

Then "\alpha=\theta=26.36\degree."


(b)

"6x^2+ x +7=6x^2-6xm+6xm+x"

"-(6m+1)(m)+(6m+1)(m)+7"

"=6x(x-m)+(6m+1)(x-m)"

"+6m^2+m+7"



"6x^2+ x +7=6x^2+12xm-12xm+x"

"-(12m-1)(2m)+(12m-1)(2m)+7"

"=6x(x+2m)-(12m-1)(x+2m)"

"+24m^2-2m+7"

The remainder is the same


"6m^2+m+7=24m^2-2m+7"

"18m^2-3m=0"

"3m(6m-1)=0"

"m_1=0, m_2=\dfrac{1}{6}"